OpenStudy (anonymous):

suppose you buy a piece of office equipment for $9,000. After 6 years you sell it for a scrap value of $5000. The equipment isdeprecated linearly over 6 years. The value of the piece of equipment ater 4 years is (round to the nearest whole dollar.

OpenStudy (anonymous):

line through \((0,9,000)\) and \((6,5000)\) is \[y=-\frac{2000}{3}x+9000\]

OpenStudy (anonymous):

put in \(4\) for \(x\)

OpenStudy (anonymous):

can you please explain this to me? I got 5000 - 9000 / 6 = -4000/6 but I cant figure out my answer to that ... which is supposed to be my M because I need to use v(t)=mt+b I thought to find the value of the piece of equipment after 4 years

OpenStudy (anonymous):

\[\frac{-4000}{6}=-\frac{2000}{3}\] same thing

OpenStudy (anonymous):

that is your \(m\)

OpenStudy (anonymous):

aka slope y intercept is what you start with \(9,000\) so line is \[y=-\frac{2000}{3}t+90000\]

OpenStudy (anonymous):

put in 4 for x

OpenStudy (anonymous):

or t

OpenStudy (anonymous):

Oh okay! How would I plug that into the TI to calculate it to find my answer?

OpenStudy (anonymous):

idk i haven't used a calculator in years http://www.wolframalpha.com/input/?i=-2000*4%2F3%2B90000

OpenStudy (anonymous):

what do you get as an answer? 87333.3333 ? what would that be to the nearest whole dollar?

OpenStudy (anonymous):

i guess \(\$87333\)

OpenStudy (anonymous):

seems wrong though hold on

OpenStudy (anonymous):

yeah it was wrong, i put in too many zeros

OpenStudy (anonymous):

\[\$6,333\] looks a lot better

OpenStudy (anonymous):

it does. why did you type it in as -2000*4/3+9000 when the 3 should be under the -2000 and multiply that by 4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it doesnt make any difference \[-\frac{2000}{3}\times 4=-\frac{2000\times 4}{3}\]